### v1

@@ 5 Conclusion
The selection of the Northern Ireland Assembly Executive Committee was based on several pre-determined agreements as well as the sequential use of D’Hondt’s method. Nonetheless, it provides a good example with which to consider the impact of sequences drawn from other apportionment methods. For instance, if we apply Theorem 21 to the top four parties who received seats, SF, DUP, AP and UUP with vote distribution $\mathbf{p}=(27,25,17,9)$, there are a total of 25 possible different sequences arising from stationary apportionments. Of course, these have length $27+25+17+9=78$. If we are interested only in the first 10 positions, we can compare sequences directly. Assuming that all parties start with no seats automatically granted, under D’Hond’ts method, the first 10 terms of the sequence are: $1,2,3,1,2,1,4,3,2,1$ (where we have used the tie break to award seat 6 to party 1).
Under Adams’ method, (applying Theorem 1), the first 10 terms of the sequence are: $1,2,3,4,1,2,3,1,2,1$, where the tie is broken in favor of SF in awarding the last seat. Both of these methods begin $1,2,3$ and both award the four parties $4,3,2$ and $1$ seats respectively (as occurred in real life); hence this will be true of all divisor methods. However, the two sequences would have very different political consequences if used to select executive committee positions one at a time. Neither of these sequences match what actually occurred, which was equivalent to the sequence $1,2,3,1,2,1,2,4,1,3$, although of course, the first three positions of First Minister, Vice Minister, and Justice Minister match what the apportionment sequences suggest. It would be interesting to investigate the consequences of different apportionment sequences in other political contexts.

### v2

@@ 5 Conclusion
The selection of the Northern Ireland Assembly Executive Committee was based on several pre-determined agreements as well as the sequential use of D’Hondt’s method. We can use this example as a way to synthesize our results and see how the sequences used to award minister positions would have changed under different stationary methods.
Referring back to Example 1, the top four parties who received ministerial positions were SF (Sinn Féin), DUP (Democratic Unionist Party), AP (Alliance Party), and UUP (Ulster Unionist Party) with vote distribution $\mathbf{p}=(27,25,17,9)$. Applying Theorem 8 to this data, there are 25 possible different sequences arising from stationary apportionment methods. Of course, these have length $27+25+17+9=78$. If we are interested only in the first 8 positions (because SF and DUP nominated the First and deputy First Ministers, respectively), we can compare sequences directly. Under D’Hondt’s method, the first 8 terms of the sequence are: $1,2,3,1,2,1,4,3$ (where we have used the tie break to award seat 6 to party 1). Under Adams’ method, applying Theorem 1, the first 8 terms of the sequence are: $1,2,3,4,1,2,3,1$. Both of these methods begin $1,2,3$ and both award the four parties $3,2,2,$ and $1$ seats, respectively (as occurred in real life); hence, this will be true if any stationary divisor method is applied (and any nonstationary divisor method, too). However, the two sequences could have different political consequences if used to select executive committee positions one at a time. The sequences for all other stationary divisor methods would be lexicographically between the sequences from the Adams and the D’Hondt methods.
Note, because AP received the Justice Minister position and their first position in the sequence was skipped (which we denote by parentheses), the actual sequence used in Northern Ireland was $1,2,(3),1,2,4,1,3$. This differs from above because their tie-break awarded UUP seat 6 instead of awarding it to SF. It would be interesting to investigate the consequences of using the apportionment sequence to also award the First and deputy First Minister positions.
@@
Appendix
Proposition 5 Let $p_{1}$ and $p_{2}$ be relatively prime with $p_{1}>p_{2}$. Assume an apportionment sequence of length $p_{1}+p_{2}$ has the form
$1^{k_{1}},\,2,\,1^{k_{2}},\,2,\,\ldots\,,1^{k_{j}},\,2,\,\ldots,\,1^{k_{p_{2}}},\,2,\,1^{p_{1}-(k_{1}+\cdots+k_{p_{2}})}.$
  1. 1.
    The apportionment sequence is generated by a stationary divisor method if and only if
    $1\leq k_{1}\leq\left\lfloor\frac{p_{1}}{p_{2}}\right\rfloor$ (1)
    and
  2. 2.
    If these conditions from inequalities (1) and (2) are met, then the cutpoint $c$ lies in the interval
    Moreover, if $\delta_{c}=\frac{(p_{1}-p_{2})}{p_{2}}c$, then
    $k_{1}$ $=\lfloor\delta_{c}\rfloor+1,$ (4)
    $k_{i}$ $=\left\lfloor\frac{p_{1}}{p_{2}}(i-1)+\delta_{c}\right\rfloor-\left\lfloor\frac{p_{1}}{p_{2}}(i-2)+\delta_{c}\right\rfloor\quad\mbox{ for }i=2,\ldots,p_{2},\text{ and}$
    $k_{p_{2}+1}$ $=p_{1}-\left\lfloor\frac{p_{1}}{p_{2}}(p_{2}-1)+\delta_{c}\right\rfloor-1.$
Proof.
Proof of part 1. To show that (1) and (2) are necessary, suppose that the sequence is generated by a stationary divisor method with cutpoint $c$. From the proof of part 2, recall this implies
or
Substituting $i=1$ into Equation (6), we get
$\frac{p_{2}(k_{1}-1)}{p_{1}-p_{2}}\leq c<\frac{p_{2}k_{1}}{p_{1}-p_{2}}.$
Because $c\in[0,1]$, this implies $\frac{p_{2}(k_{1}-1)}{p_{1}-p_{2}}<1$ and $\frac{p_{2}(k_{1})}{p_{1}-p_{2}}>0$. Solving for $k_{1}$ and recalling that $k_{1}$ is an integer, it follows that $1\leq k_{1}\leq\left\lfloor\frac{p_{1}}{p_{2}}\right\rfloor$.
To show that Equation (2) is necessary, note that in order for $c$ to exist, the intervals in Equation (6) must have nonempty intersection. Thus, each left end-point must be to the left of each right end-point. So for each $i\neq i^{\prime},$
If $i<i^{\prime}$, this implies $p_{1}(i^{\prime}-i)<p_{2}(k_{i+1}+\cdots+k_{i^{\prime}}+1)$. If $i^{\prime}<i$ then $p_{1}(i-i^{\prime})>p_{2}(k_{i^{\prime}+1}+\cdots+k_{i}-1)$. Combining these (and switching the labels $i$ and $i^{\prime}$ in the second inequality so that $i^{\prime}>i$), we get
or
Equation (2) follows because the $k_{i}$ are integers.
Conversely, suppose that the apportionment sequence has the given form and that Equations (1) and (2) are satisfied. Then the interval in Equation (3) is non-empty and if $c$ lies in the interval, it clearly satisfies Equation (5) for each $i$. Hence the stationary divisor method with cutpoint $c$ yields the same sequence. It remains to show that the interval (3) lies in $[0,1].$ To see this, note that for each $i\geq 1$,
Thus, the right-hand side of the interval from Equation (3) is greater than 0. Similarly, we can show the left-hand side of the interval (3) is less than or equal to 1, implying that $c\in[0,1]$. ∎
Theorem 3 Suppose $p_{1}$ and $p_{2}$ are relatively prime and that $p_{1}=ap_{2}+b$ for some integers $a\geq 1$ and $0<b<p_{2}$. If $\delta_{c}=\frac{(p_{1}-p_{2})}{p_{2}}c=(a-1+\frac{b}{p_{2}})c$, then
$k_{1}$ $=\lfloor\delta_{c}\rfloor+1,$
$k_{i}$ $=\left\{\begin{array}[]{cl}a+1&\mbox{ if }i\in S\\ a&\mbox{ if }i\notin S\end{array}\right.,\mbox{ and}$
$k_{p_{2}+1}$ $=a+b-1-\left\lfloor\frac{b}{p_{2}}(p_{2}-1)+\delta_{c}\right\rfloor$
where
$S=\left\{i=2,\ldots,p_{2}:i=\left\lceil(j-\delta_{c}+\left\lfloor\delta_{c}\right\rfloor)\frac{p_{2}}{b}\right\rceil+1\mbox{, for }j=1,2,\ldots,b\right\}.$
Proof.
The formulas for $k_{1}$ and $k_{p_{2}+1}$ follow directly by substituting $p_{1}=ap_{2}+b$ into Equation (4). Doing the same for the remaining $k_{i}$, we obtain for $i=2,\ldots,p_{2}$,
$k_{i}$ $=\left\lfloor\frac{p_{1}}{p_{2}}(i-1)+\delta_{c}\right\rfloor-\left\lfloor\frac{p_{1}}{p_{2}}(i-2)+\delta_{c}\right\rfloor$
$=\left\lfloor a(i-1)+\frac{b}{p_{2}}(i-1)+\delta_{c}\right\rfloor-\left\lfloor a(i-2)+\frac{b}{p_{2}}(i-2)+\delta_{c}\right\rfloor$
$=a+\left\lfloor\frac{b}{p_{2}}(i-1)+\delta_{c}\right\rfloor-\left\lfloor\frac{b}{p_{2}}(i-2)+\delta_{c}\right\rfloor.$
Since $\frac{b}{p_{2}}<1$, these exponents are equal to either $a$ or $a+1$. We claim at most $b$ of the $k_{i}$ are equal to $a+1$. To see this, we substitute $i^{\prime}=p_{2}$, $i=1$ and $p_{1}=ap_{2}+b$ into Equation (2), to obtain
Finally, to determine which of the $k_{i}$ is equal to $a+1$, consider the quantity $K_{i}=\lfloor(i-1)\frac{b}{p_{2}}+\delta_{c}\rfloor$. If $i=1$, $K_{i}=\lfloor\delta_{c}\rfloor$. As $i$ increases, the values of $K_{i}$ are constant until some smallest $i$ (call it $i_{1}$) satisfies $K_{i_{1}}=\lfloor\delta_{c}\rfloor+1$; thus $i_{1}$ is the smallest index such that $k_{i}=a+1.$ Similarly, if $i_{2}$ is the smallest $i$ such that $K_{i_{2}}=\lfloor\delta_{c}\rfloor+2$ then $i_{2}$ is the second smallest index such that $k_{i}=a+1.$ More generally, the indices for which $k_{i}=a+1$ occur at the smallest indices $i_{j}$ for which $K_{i_{j}}=\lfloor\delta_{c}\rfloor+j.$
Let $\ell_{j}=\lceil(j-\delta_{c}+\lfloor\delta_{c}\rfloor)\frac{p_{2}}{b}\rceil+1$; we claim $i_{j}=\ell_{j}$ for $j\geq 1$. To see this, note that
$(j-\delta_{c}+\lfloor\delta_{c}\rfloor)\frac{p_{2}}{b}\leq\ell_{j}-1<(j-\delta_{c}+\lfloor\delta_{c}\rfloor)\frac{p_{2}}{b}+1,$
which implies
$j+\lfloor\delta_{c}\rfloor\leq(\ell_{j}-1)\frac{b}{p_{2}}+\delta_{c}<j+\lfloor\delta_{c}\rfloor+\frac{b}{p_{2}},$
and hence $\lfloor(\ell_{j}-1)\frac{b}{p_{2}}+\delta_{c}\rfloor=j+\lfloor\delta_{c}\rfloor.$
On the other hand
$(j-\delta_{c}+\lfloor\delta_{c}\rfloor)\frac{p_{2}}{b}-1\leq\ell_{j}-2<(j-\delta_{c}+\lfloor\delta_{c}\rfloor)\frac{p_{2}}{b},$
which implies
$j+\lfloor\delta_{c}\rfloor-\frac{b}{p_{2}}\leq(\ell_{j}-2)\frac{b}{p_{2}}+\delta_{c}<j+\lfloor\delta_{c}\rfloor,$
and so $\lfloor(\ell_{j}-2)\frac{b}{p_{2}}+\delta_{c}\rfloor<j+\lfloor\delta_{c}\rfloor.$ Thus, $\ell_{j}$ is the smallest integer such that $K_{\ell_{j}}=\lfloor\delta_{c}\rfloor+j$ or $i_{j}=\ell_{j}$, as required. ∎
Theorem 4 Suppose $p_{1}$ and $p_{2}$ are relatively prime and $p_{1}>p_{2}+1$. Then the apportionment sequences corresponding to stationary divisor methods are constant for cutpoints in the intervals $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ where $\ell=0,\ldots,p_{1}-p_{2}-1$. Moreover, each of these $p_{1}-p_{2}$ sequences is distinct.
Proof.
Suppose $p_{1}=ap_{2}+b$ for some integers $a\geq 1$ and $0<b<p_{2}$ and suppose that $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ for some $\ell=0,\ldots,p_{1}-p_{2}-1$. Let $\ell=\nu p_{2}+\ell^{\prime}$ for some integers $\nu\geq 0$ and $0\leq\ell^{\prime}<p_{2}.$ Then $\delta_{c}=\frac{p_{1}-p_{2}}{p_{2}}c$ satisfies $\frac{\ell}{p_{2}}\leq\delta_{c}<\frac{\ell+1}{p_{2}}$, which implies $\lfloor\delta_{c}\rfloor=\nu$ and $k_{1}=\nu+1$. Thus, $k_{1}$ is constant on this interval.
To show that the other $k_{i}$ are constant on $\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ it suffices to show that the indices $i_{j}=\lceil(j-\delta_{c}+\lfloor\delta_{c}\rfloor)\frac{p_{2}}{b}\rceil+1$, from the definition of $S$ in Theorem 3, are also constant on this interval. But
$\frac{-\ell^{\prime}-1}{p_{2}}<-\delta_{c}+\lfloor\delta_{c}\rfloor\leq\frac{-\ell^{\prime}}{p_{2}},$
which implies
$\frac{jp_{2}-\ell^{\prime}-1}{b}<(j-\delta_{c}+\lfloor\delta_{c}\rfloor)\frac{p_{2}}{b}\leq\frac{jp_{2}-\ell^{\prime}}{b}.$
So $i_{j}=\lceil(j+\lfloor\delta_{c}\rfloor-\delta_{c})\frac{p_{2}}{b}\rceil+1=\lceil\frac{jp_{2}-\ell^{\prime}}{b}\rceil+1$ is constant on the interval as claimed.
Suppose that $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ and $c^{\prime}\in[\frac{\ell+1}{p_{1}-p_{2}},\frac{\ell+2}{p_{1}-p_{2}})$ and that the two apportionment sequences are the same. Then $\delta_{c^{\prime}}=\frac{p_{1}-p_{2}}{p_{2}}c^{\prime}$ satisfies $\frac{\ell+1}{p_{2}}\leq\delta_{c^{\prime}}<\frac{\ell+2}{p_{2}}$, and recalling that $\ell=\nu p_{2}+\ell^{\prime},$ we obtain $\nu+\frac{\ell^{\prime}+1}{p_{2}}\leq\delta_{c^{\prime}}<\nu+\frac{\ell^{\prime}+2}{p_{2}}.$ Since the $k_{1}$ values for each sequence are equal, $\lfloor\delta_{c^{\prime}}\rfloor=\lfloor\delta_{c}\rfloor=\nu$. Hence $\ell^{\prime}+1<p_{2}$.
$\left\lceil\frac{jp_{2}-\ell^{\prime}}{b}\right\rceil=\left\lceil(j-\delta_{c}+\lfloor\delta_{c}\rfloor)\frac{p_{2}}{b}\right\rceil=\left\lceil(j-\delta_{c^{\prime}}+\lfloor\delta_{c^{\prime}}\rfloor)\frac{p_{2}}{b}\right\rceil=\left\lceil\frac{jp_{2}-(\ell^{\prime}+1)}{b}\right\rceil.$